# XXLTraffic: Expanding and Extremely Long Traffic Dataset for Ultra-Dynamic Forecasting Challenges

### v1

@@ A.1 Dataset
Table 6: Overview of Gap, Hourly, and Daily Aggregated Data Based on Processed Raw Data
@@ A.2 Results
Table 7: Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol -- indicates that the result is an outlier.
@@
Table 8: Comparison in gap dataset in District 05,06,10,11,12. The bold text indicates the best.
@@
Additionally, we provided the results of naive zero-shot forecasting, as shown in Table 9. The poor performance of this method indicates significant potential for improvement.
Table 9: Results of zero shot forecasting in PEMS03. In this study, we utilized the inputs from 151 existing nodes in PEMS03 to predict 52 new nodes, resulting in a test set comprising a total of 203 nodes.
Datasets 96 192 336
Metrics MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year gap 4.658 3.465 4.945 3.648 5.198 3.892
1.5-year gap 4.891 3.657 5.124 3.842 5.263 3.996
2-year gap 4.547 3.410 4.895 3.539 5.103 3.758

### v2

@@ A.1. Dataset
Table 6. Overview of Gap, Hourly, and Daily Aggregated Data Based on Processed Raw Data
@@ A.2. Results
Table 7. Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol -- indicates that the result is an outlier.
@@
Table 8. Comparison in gap dataset in District 05,06,07,10,11,12. The bold text indicates the best.
@@
[…]
Table 9. Model Training Time Comparison. The training time for all baselines per epoch is measured in seconds.
Baselines Mamba iTransformer DLinear Autoformer Informer FEDFormer MICN PatchTST STGCN ASTGCN GWN AGCRN PDFormer
PEMS04_gap1 396.3 1310.2 237.9 1656.7 652.5 900.9 409.8 3649.4 2425.9 7921.9 4110.4 13369.1 8013.9
tfNSW_gap1 13.7 44.1 7.4 148.2 73.3 380.1 101.2 55.3 37.3 228.8 50.4 658.6 47.9